OpenStudy (anonymous):

find lim as x goes to 0 2/(x csc x)

OpenStudy (anonymous):

does it help to know that \[\frac{2}{\csc(x)}=2\sin(x)\]?

OpenStudy (aum):

\[\large \lim_{x \rightarrow 0}\frac{2}{x\csc(x)} = \lim_{x \rightarrow 0}\frac{2\sin(x)}{x} = 2 * \lim_{x \rightarrow 0}\frac{\sin(x)}{x} = ? \]